1 2 8 48 384
100 96 104 88 120 56
A Z E B I Y O
What comes after each series?

Answer by Shai(25)

Thats pretty easy,
100 96 104 88 120 56
If one would closely observe
(100-96)=-4
(96-104)=8
(104-88)=-16
(88-120)=32
(120-56)=-64
The series here has
1)alternate positive and negative signs
2)also,twice (negative)the first number gives the
Subsequent number i.e (-4*-2)=8 (8*-2)=-16 and so
On
If you look closely the number after 56
Would be 56+(-64*-2)=184